I need help solving each system of equations. Here are the two question:
x+y-z=4
y =6
y+z=13

Answer by richwmiller(17219)

One problem per submission.
THe first one is very easy
y =6
y+z=13
x+y-z=4
take them in this order
y=6
substitute 6 for y in the 2nd
6+z=13
z=7
now sub in the third
x+6-13=4
add 13
x+6=17
subtract 6
x=11